OpenStudy (anonymous):

Solve for the variable : -x/9 = 5/3

jimthompson5910 (jim_thompson5910):

Cross multiply to get... \[\Large \frac{-x}{9} = \frac{5}{3}\] \[\Large -x * 3 = 9*5\] \[\Large -3x = 45\] I'll let you take over from here

OpenStudy (anonymous):

x = -15

jimthompson5910 (jim_thompson5910):

good you got it